OpenStudy (anonymous):

An 18-foot pole is leaning against a wall. It touches the wall 16 feet above the ground. Find the measure of the angle that the pole makes with the ground. Round your answer to the nearest degree.

OpenStudy (anonymous):

\[180-90-cos^{-1}(\frac{16}{18})\] As there's 180 degrees in a triangle, and the wall is probably at a right angle to the ground.

OpenStudy (anonymous):

Actually, you can just do: \[sin^{-1}(\frac{16}{18})\] Sorry, haven't done trig in a while..

OpenStudy (anonymous):

sinθ = 16/18 = 8/9 θ = arcsin(8/9) = 62.73° θ = 63°

OpenStudy (anonymous):

\[\sin^{-1}(\frac{8}{9})\] what stroodle said

OpenStudy (anonymous):

also what mathematicalpro said 62.7339555...